OpenStudy (sarahc):

HELP PLEASE! Graph the ellipse with equation x squared divided by 9 plus y squared divided by 25 = 1.

OpenStudy (anonymous):

\[\frac{x^2}{9}+\frac{y^2}{25}=1\]?

OpenStudy (sarahc):

yes

OpenStudy (anonymous):

we have a choice we can do it step by step or go right to the answer do you need to know how to graph this one your own?

OpenStudy (sarahc):

right to the answer is fine

OpenStudy (sarahc):

@satellite73

OpenStudy (sarahc):

@aaronq @jim_thompson5910